OpenStudy (anonymous):

Eight less than a number doubled is twenty more than the number. What is the number?

OpenStudy (anonymous):

X*2-8=20+x Add 8 to both sides 2x=28+x Subtract x from both sides 1x=28 X=28

OpenStudy (anonymous):

8 less than "a number" can be represented as x (the number) -8. 8 less than "a number doubled" would be 2*x-8. So 2x-8 equals x (the number) plus 20 more, or 20+x. Isolate x, and it equals 28

OpenStudy (anonymous):

thank you very much